The probability that the birthdays of five persons chosen at random will fall in twelve different calender months is zero.
You would need at least twelve persons to have a non zero probability.
5/20, 25%, or .25
Describe an event that has a probability of 1 and an event that has a probability of 0?
P(a^~a)==1 P(a&~a)==0
the line above is shorthand notation for an event that has a probability of 1, followed by an event that has a probability of 0. P(event) is an easy way to say the probability of "event". The "^" means "OR", the "~" means "NOT", and the "&" as you are probably familiar means "AND". So puting it together, "P(a^~a)" means the probability that an event "a" occurs OR that event "a" does not occur. So take an event "a", any event, like drawing an 8 of clubs out of a deck of cards. If you draw a random card out of a deck of cards, the probability that you will draw an 8 of clubs OR that you will not draw an 8 of clubs is 1. That means 100%. So when you draw a card you will either draw the 8 of clubs or not draw the 8 of clubs. It seems like an obvious statement to make but its a proof that becomes very important in proving less obvious theories. likewise, the second statement was "P(a&~a)", so the probability that event "a" occurs AND event "a" doesnt occur. Since the event has to either occur or not, it cant occur AND not occur, so the probability is 0.
Probability to get exactly 2 heads when flipping a coin 3 times?
The probability of throwing exactly 2 heads in three flips of a coin is 3 in 8, or 0.375.
There are 8 outcomes of flipping a coin 3 times, HHH, HHT, HTH, HTT, THH, THT, TTH, and TTT. Of those outcomes, 3 contain two heads, so the answer is 3 in 8.
First you need to figure out the probability of it not landing on a 2 when you roll one die. Since there are six sides and two sides with 2 on it, the probability that one die will not land on 2 is 4/6.
Now you multiply 4/6 by 4/6 to find the correct odds as to both die not landing on 2.
4/6 * 4/6 = 16/36 = 4/9
The percentage is (4/9)*100% = (approx.) 44.4%
Prob(Sum is odd) = 1/2
Prob(Sum > 6) = 15/36 = 5/12
Prob(Sum is odd or >6) = 27/36 = 3/4
What does the standard deviation tell us about the mean?
The standard deviation tells us nothing about the mean.
If you roll the die one time, there is a 1/3 probability that you will get a 1 or a 3.
There are 6 possibilities: 1,2,3,4,5,6-----> This number goes on the bottom of your fraction because it is what the probability is out of.
You want 2 different numbers: 1,6--------> This goes on the top because those are the 2 numbers you want.
So the fraction (probability) is 2/6, which simplifies to 1/3.
Probability of getting all questions wrong on multiple choice test?
In order to answer, the number of questions on the test must be given.
Using mean of 6.375 and Standard deviation of 1.47 plot values on normal distribution?
Your middle point or line for the plot (mean) would be 6.375. Then you would add/subtract 1.47 from your mean. For example, one standard deviation would equal
6.375 + 1.47 and one standard deviation from the left would be
6.375 - 1.47
What are the Three Approaches to finding a Probability?
1. subjective probability (intelligent guess)
2. relative frequency (in percent)
3. classical probability (in decimal)
What is the probability that you will spin a 4 and roll a 4?
Assuming each possible number on a spinner has the same probability and an unbiased die is being rolled, the answer depends on how many numbers are on the spinner, and how many times the number 4 appears on each.
To find the probability, workout the probability of spinning a 4 on the spinner and the probability of rolling a 4 on the die; then as spinning the spinner has no effect on rolling the die, they are independent events and to get the probability of both happening multiply them together.
The probability of success is the number of successful outcomes divided by the total number of outcomes, giving:
Probability(spinning a 4) = how_many_4s_are_on_the_spinner / how_many_numbers_are_on_the_spinner
Probability(rolling a 4) = how_many_4s_are_on_the_die / how_many_numbers_are_on_the_die
Probability(spinning a 4 and rolling a 4) = Probability(spinning a 4) × Probability(rolling a 4)
Examples:
Probability values are never negative and are always between 0-1 according to the definition Probability of A= Number of outcomes classified as A/Total number of possible outcomes
Why do population follow normal distribution?
The form of this question incorportates a false premise. The premise is that the data are normally distributed. Actually, is the sample mean which, under certain circumstances, is normally distributed.
If a family has three children find the probability that two of the children are girls?
there is a 50% chance that two of them will be girls
What is the exponential growth formula?
y = a * bxUsing "^" for powers. Any exponential formula can be written in one of the following two forms:
y = A e^Bx
y = C D^x
where "y" is the independent variable, "x" is the dependent variable, and "A", "B", "C", and "D" are arbitrary constants.
Find the probability that the month of February in a year 2008 may have 5 Tuesdays?
Considering the fact that a non-leap year having 28days in February will definitely have 4 Tuesdays , we have to find the probability that the 29th day of a leap year (2008) is also a Tuesday. This probability is 1/7 i.e 1 out the probable 7days in a week !
Well this said, this question is void in a sense knowing that we are in 2010 and that February 2008 had 5 Fridays and therefore the probability of 5 Tuesdays is ZERO as of 2008 !